# Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction

### v4

@@ 6.4 Avenues for Future Research
    Richer interpretability workflows: integrate SHAP-on-decoder analyses [62], textual factor labeling from analyst reports, and practitioner validation workshops.
@@
    Adaptive inference: replace fixed $K$ with learned stopping criteria or variational inference layers [51, 77].
@@
    Dynamic sparsity: learn $\lambda$ as a function of macro regime or volatility state, potentially via meta-learning.
@@
    Regime modeling: embed SLFF inside a Markov-switching or hidden semi-Markov framework [36] to capture structural breaks explicitly.
@@
    Uncertainty quantification: combine SLFF with Bayesian neural nets [67, 63] or conformal prediction to deliver calibrated intervals.
@@
    Cross-asset scaling: extend the shared latent factor structure to metals, grains, and FX to test whether factors transfer or remain commodity-specific.
@@
    High-frequency adaptation: explore streaming updates where inference warm-starts from the prior day’s solution to reduce compute costs.
@@
    Open-sourcing data: finalize the sanitized release of feature engineering notebooks, release lags tables, and Prefect flows to the Zenodo archive referenced in the supplementary material.
  9. 9.
    Broader benchmarking: add diffusion-based sequence models (e.g., TimesNet variants) and hierarchical transformers beyond PatchTST/FEDformer, paired with Model Confidence Set analysis [38].
@@ 6.5 Conclusion
By reframing the method as a sparse latent factor forecaster with iterative inference, we bridge the gap between training-time optimization and test-time deployment, enforce information-set discipline, and benchmark against state-of-the-art sequence models under rolling-origin evaluation. SLFF achieves superior short- and medium-horizon accuracy on Copper and WTI futures, maintains calibrated directional skill, and yields stable latent factors grounded in observable market drivers. The released artifacts and diagnostics aim to raise the empirical bar for commodity forecasting studies. Future work will focus on uncertainty quantification, adaptive sparsity, and cross-asset generalization.
Appendix A Diagnostic Configuration and Sensitivity Checks
This appendix summarizes diagnostic results that use a reduced configuration ($m=3$ with normalized targets) to study inference alignment and sensitivity. These diagnostics are not part of the main performance tables.
Path RMSE MAE
Refined ($\mathbf{z}^{\star}\rightarrow\text{Dec}$) 0.5498 0.4294
Refinement gain ($\Delta$) 0.1428 0.0973
Sensitivity checks. Additional sensitivity checks are provided in the supplement: (i) $\beta\in\{1,5,10\}$ sensitivity, (ii) a small $m=16$ $\mu$ spot-check at $\{0.05,0.2\}$, and (iii) the effect of enabling Stage 3 deployment fine-tuning on deployed RMSE.
Appendix B Public-Only Proxy Mapping and Driver Exclusions
Table 14 lists the public-only proxy mapping used when proprietary feeds are unavailable. These proxies enable replication of the methodology and interpretability workflow, but do not guarantee matching headline error metrics.
Table 14: Public-only proxy mapping for proprietary feeds.
Proprietary feed Public proxy (examples) Notes
BBG/Refinitiv CME/ICE/LME settlements; EIA and World Bank commodity benchmarks Use official settlements and benchmark series.
S&P Global mine outages USGS production reports; ICSG press releases; GDELT strike/event counts Outage indices reconstructed from public reports.
Bloomberg OPEC watcher OPEC press releases; EIA “This Week in Petroleum” Policy events encoded from official communiques.
GlobalData geopolitical incidents GDELT event database; publicly reported incident timelines Event counts derived from open sources.
Held-out drivers (not in $X_{t}$). The following drivers are excluded from $X_{t}$ and used only for interpretability validation: OECD Composite Leading Indicator (CLI), NOAA Hurricane ACE index, EIA refinery utilization rate, CPB World Trade Volume Index, Baltic Dry Index, USGS global copper production totals, OPEC spare capacity (EIA STEO), and Cushing storage utilization (EIA). Overlapping drivers (e.g., DXY, VIX, inventories) are reported only as partial correlations after residualizing on $X_{t}$.
Appendix C Latent Factor Correlation Analysis
This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in (Section 5.5). We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes). These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations. It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
C.1 Constructed Event Indices
@@
C.2 Copper Latent Factor Correlations
@@
  • Factor $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory Dynamics): This factor shows a strong positive correlation (+0.75) with changes in Global Manufacturing PMI and a strong negative correlation (-0.65) with changes in LME Copper Inventories. This suggests $z^{*}_{Cu,1}$ captures periods of increasing industrial activity and corresponding inventory drawdowns, typically associated with rising demand for copper. Its weaker negative correlation with the USD Index (-0.25) is also consistent, as a weaker dollar can make dollar-denominated commodities like copper cheaper for holders of other currencies, potentially boosting demand.
  • Factor $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD Influence): This factor is strongly positively correlated (+0.70) with changes in the USD Index (DXY). It also shows a moderate positive correlation (+0.35) with the VIX. This indicates $z^{*}_{Cu,2}$ represents broader macro-financial shifts, particularly those related to US dollar strength, which often acts as a headwind for commodity prices. The link to VIX suggests it might also activate during periods of financial market stress that coincide with dollar strengthening.
  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with the observed Supply Disruption Index (Appendix C), where lower values denote more severe outages. This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values. Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX, suggesting it captures periods of heightened market fear or risk aversion. Its moderate positive correlation with the DXY change (+0.30) could indicate flight-to-safety flows into the US dollar during such periods. This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
C.3 Crude Oil Latent Factor Correlations
@@
  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with DOE product supplied (gasoline demand). This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with the Geopolitical/OPEC+ event index, where positive values signal supply-tightening announcements. It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility. This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility, which can be associated with broader risk-off sentiment or increased speculative activity. Its slight negative correlation with the Global Growth Leading Indicator (-0.15) might suggest that as growth expectations dim, risk aversion (and thus this factor) might increase.
  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator. This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market. Its moderate positive correlation with the DOE gasoline product supplied series (+0.20) is also consistent with general demand trends.
These interpretations, based on correlations, provide a plausible understanding of how the SLFF model might be decomposing market information into sparse, meaningful drivers. Further validation using more formal attribution methods (as discussed in Section 5.5) would be required to confirm these relationships more rigorously.
Appendix D Event Studies and Counterfactuals
This appendix lists the event windows (pre/post days), factor responses, and forecast impacts used in the counterfactual study. Each table includes the event definition (e.g., OPEC+ surprise cuts, ICSG mine shutdown announcements, CFTC positioning extremes), the latent factors most affected, the t-statistics of the factor change relative to a randomization null, and the induced shift in 5-day and 22-day forecasts. Scripts to regenerate these tables are part of the released interpretability notebooks.

### v5

@@ 8.7 Future Research Directions
    Cross-commodity sensitivity schedules: Move beyond fixed $\mu=0.1$ by learning adaptive $(\mu,\lambda)$ schedules that respond to volatility and regime changes.
@@
    Causal factor validation: Apply causal discovery methods to validate whether factor correlations reflect causal influence. Collaborate with domain experts.
@@
    Adaptive sparsity: Learn $\lambda$ as a state-dependent function via meta-learning or curriculum learning.
@@
    Multivariate targets: Extend decoder to jointly predict log-prices, spreads, and implied volatility.
@@
    Cross-asset latent sharing: Train SLFF on four commodities (Copper, Crude Oil, Natural Gas, Gold) with shared sparse latents plus commodity-specific decoders.
@@
    Uncertainty quantification: Develop Bayesian variant or integrate conformal prediction for calibrated intervals.
@@
    Streaming/online updates: Implement warm-start inference where day-$t$ loop initializes at day-t1t-1 refined solution, reducing compute.
@@
    Broader benchmarking: Compare against diffusion-based forecasters and hierarchical transformers using rolling-origin protocol and Model Confidence Set analysis.
@@ 8.8 Conclusion
We have introduced the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a method that explicitly addresses the amortization gap in latent-variable forecasting via sparse iterative refinement and encoder alignment training. We ground the method in design intuitions derived under a linearized decoder (amortization gap bounds and convergence rates, empirically validated against the deployed MLP), implement rigorous information-set controls to prevent mixed-frequency data leakage, and formalize a three-stage interpretability protocol. On commodity futures (Copper, WTI, Gold; 2013–2025), SLFF achieves statistically significant forecast improvements at 1- and 5-day horizons (DM $p<0.035$) and yields sparse latent factors that are stable across seeds and correlated with observable economic drivers. At the 22-day horizon, SLFF and leading baselines are statistically indistinguishable ($p>0.10$), indicating that gains are concentrated where short-horizon information has not yet been fully priced.
Beyond commodities, we provide a taxonomy of problem characteristics for assessing SLFF’s applicability, identify candidate domains (electricity, macro nowcasting, agriculture) that satisfy the structural criteria but remain empirically unvalidated, and design synthetic experiments confirming that SLFF recovers ground-truth sparse factors under controlled conditions. The core methodological contribution—addressing the amortization gap via iterative refinement + encoder alignment—is applicable to any latent-variable forecasting problem. The information-set and interpretability protocols generalize to other domains. Reproducibility artifacts, code, and diagnostic notebooks are released to enable peer validation and extension.
Appendix A Latent Factor Correlation Analysis
This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in Section 6.5. We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes). These correlations are correlational and exploratory and should be interpreted carefully; they do not imply causality. Observable relationships form the basis for tentative economic interpretations. It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
A.1 Constructed Event Indices
@@
A.2 Copper Latent Factor Correlations
@@
  • Factor $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory Dynamics): This factor shows a strong positive correlation (+0.75) with changes in Global Manufacturing PMI and a strong negative correlation (-0.65) with changes in LME Copper Inventories. This is consistent with industrial demand and inventory dynamics; factors exert no claimed causal effects.
  • Factor $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD Influence): This factor is strongly positively correlated (+0.70) with changes in the USD Index (DXY). This indicates a correlational relationship with macro-financial conditions, particularly US dollar strength.
  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with the observed Supply Disruption Index. This indicates a strong correlational relationship; the interpretation is that extreme values of this factor correlate with periods of supply disruptions.
  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX. This suggests it is associated with periods of heightened market stress or uncertainty.
A.3 Crude Oil Latent Factor Correlations
@@
Figure 5: Heatmap of Correlations: Gold Latent Factors vs. Selected Features (Test Set).
@@
  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with DOE product supplied (gasoline demand). This is consistent with the state of the US oil market balance: higher values correlate with falling inventories and/or rising demand.
  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with the Geopolitical/OPEC+ event index. This suggests a strong association with major supply-side policy decisions and geopolitical disruptions.
  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility.
  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator. This indicates a correlational relationship with broader global demand expectations.
These correlations provide an exploratory understanding of how the SLFF model may decompose market information into sparse, meaningful patterns. Further validation using more formal causal attribution methods (as discussed in Section 6.5) would be required to confirm these relationships more rigorously.
Appendix B Event Studies and Counterfactuals
Event studies anchor latent factor responses to significant market and policy announcements. Each includes event dates, affected factors, test statistics, and induced forecast shifts. Randomization null: 1000 permutations.
Table 17: Aggregated Event Study: Copper Mine Disruptions (n=18, 2017-2023). Window: $\pm 3$ days. Affected: $z_{Cu,3}$. Power >0.90>0.90 for $d=1.0$.
Aggregate Shift (SD) t-stat p-value d 5/22-day Shift (%)
All Disruptions -0.79 -5.23 <<0.001 1.45 -1.0 / -2.1
Table 18: Aggregated Event Study: OPEC+ Policy Shocks (n=15, 2017-2023). Window: $\pm 3$ days. Affected: $z_{Oil,2}$. Power >0.90>0.90 for $d=1.0$.
Aggregate Shift (SD) t-stat p-value d 5/22-day Shift (%)
All OPEC+ Policy Shocks +0.88 4.97 <<0.001 1.31 +1.8 / +3.7
Shifts are coherent with economics and sufficiently powered at the aggregate level; interpretation remains correlational rather than causal.